# Spiking Neural Networks for Radio Frequency Interference Detection in Radio Astronomy

### v1

@@ Abstract
Spiking Neural Networks (SNNs) promise efficient spatio-temporal data processing owing to their dynamic nature. This paper addresses a significant challenge in radio astronomy, Radio Frequency Interference (RFI) detection, by reformulating it as a time-series segmentation task inherently suited for SNN execution. Automated RFI detection systems capable of real-time operation with minimal energy consumption are increasingly important in modern radio telescopes. We explore several spectrogram-to-spike encoding methods and network parameters, applying first-order leaky integrate-and-fire SNNs to tackle RFI detection. To enhance the contrast between RFI and background information, we introduce a divisive normalisation-inspired pre-processing step, which improves detection performance across multiple encoding strategies. Our approach achieves competitive performance on a synthetic dataset and compelling results on real data from the Low-Frequency Array (LOFAR) instrument. To our knowledge, this work is the first to train SNNs on real radio astronomy data successfully. These findings highlight the potential of SNNs for performing complex time-series tasks, paving the way for efficient, real-time processing in radio astronomy and other data-intensive fields.
@@
Spiking Neural Networks (SNNs) are inspired by how biological neurons communicate. Instead of using continuous signals, as in Artificial Neural Networks (ANNs), they transmit information in discrete spikes or pulses, closer to how natural neurons fire [1]. ANNs have proven to be highly effective, especially when deployed on conventional computing hardware platforms, and they excel in numerous tasks when combined with large datasets. The central question remains whether SNNs can reliably achieve comparable or superior results by mimicking biological neurons more closely while using less energy and computational power, particularly when running on neuromorphic computing systems [2]. The potential for SNNs to outperform traditional models hinges on developing efficient neuromorphic hardware and identifying data-intensive applications that can exploit the temporal dynamics unique to SNNs [3]. While SNNs have shown promise in several domains, particularly those leveraging neuromorphic sensing hardware, tackling increasingly demanding tasks is vital to moving the field forward.
@@
Both conventional and machine-learned approaches treat RFI detection as a two-dimensional semantic segmentation or anomaly detection problem, where spectrograms, or ‘visibilities,’ are treated like images. Although effective, existing two-dimensional approaches to RFI detection necessitate completing an entire observation before processing can commence. This is a significant limitation as data destined to be removed is stored and processed in costly high-performance computing facilities [4]. As such, there is a strong imperative in the field to develop RFI detection techniques that are accurate and capable of operating in near real-time, with minimal energy consumption, and where applicable, minimal training data [14, 17].
@@
Previous applications of SNNs to spectrographic data, such as audio processing, have primarily focused on classification tasks [21, 22]. In such tasks, a time-varying input is processed, but the final output is reduced to a single value or decision derived from the final layer of the network. However, the problem of RFI detection requires the output to retain the same temporal and spatial resolution as the input data. In this work, we reformulate RFI detection as a time-series segmentation problem, where the network must output a prediction that spans the same time and frequency dimensions as the input spectrogram to explore how well first-order leaky-integrate and fire (LiF) spiking neurons can capture temporal information in this task. To our knowledge, this is the first time SNNs have been applied to this kind of task, and until now, it has been unknown whether backpropagation through time (BPTT) can handle the complexity of such tasks.
@@
This paper provides a robust formulation of RFI as a time-series segmentation task appropriate for direct SNN execution, investigates encoding visibility data with latency, rate, delta-modulation, delta-exposure and three distinct variations on step-forward encoding, extensive hyper-parameter optimisation, a divisive normalisation inspired pre-processing step which greatly improves detection performance across several encoding methods, and results for SNNs trained and optimised with each encoding for a synthetic dataset simulating the Hydrogen Epoch of Reionisation Array (HERA) in South Africa, and a real RFI detection dataset derived from the Low-Frequency Array (LOFAR) instrument in the Netherlands with comparison to identically sized ANN feed-forward networks. This paper builds upon our previously presented work [25], extending it in several important ways. The delta-exposure encoding method and divisive-normalisation-inspired pre-processing step are previously unseen techniques. We provide more comprehensive hyper-parameter optimisation extending into network depth and width in addition to a more extensive analysis of the resulting parameter space. Lastly, by including results for the real LOFAR dataset, we present the first study demonstrating the application of from-scratch trained SNNs on real radio astronomy data. In doing so, we achieve state-of-the-art RFI detection performance on the synthetic HERA dataset and competitive results on the real LOFAR dataset, demonstrating the efficacy of our approach while employing small multi-layer feed-forward SNNs. This investigation demonstrates the capability of SNNs to perform a data-intensive complex time-series segmentation task in the form of RFI detection with relatively simple network architectures, paving the way for more sophisticated SNN-based RFI detection schemes and critically finding a scientifically impactful problem in which SNNs and neuromorphic computing may hold unique advantages over contemporary methods.
@@ Results
RFI Detection as Time-Series Segmentaton
@@
This formulation tasks a classifier to segment the time series, in this case, as boolean values rather than a two-dimensional image. In the case of an ANN, the encoding and decoding functions are the identity. This formulation leaves encoding and decoding visibility data as spikes for this specific task open for experimentation. The spike encoding process involves stretching the original time steps in the spectrograms into longer exposures, allowing the SNN to react to the current input. We experimented with seven encoding methods, which we briefly outline. The methods section presents detailed formulations and examples of each encoding approach.
Latency
Latency encoding transforms pixel intensity into spike timings. This method benefits from a deterministic decoding method and a fixed number of spikes.
Rate
Rate encoding translates pixel intensities into spike rates, interpreting intensity as the probability of a Poisson process for each exposure time. Rate decoding is inherently more stochastic than most other methods, so achieving reliable results in this task is difficult. Moreover, spike counts increase proportionally to the exposure time.
Delta-Modulation
Delta-modulation encoding fires a positive polarity spike when a positive difference between two time steps passes a threshold and a negative polarity spike when a negative difference passes a threshold. Delta-modulation is easily decoded but can struggle with frequency-localised RFI.
Delta-Exposure
Delta-exposure encoding is similar to delta-modulation but includes time for the network to react to each time step in the original spectrogram.
Step-Forward
Step-forward encoding comes from Speech2Spikes [21] and is inspired by encoding vocal spectrograms. This method encodes the input’s transient and cumulative properties in two separate sections, doubling the input-width. We experimented with three sub-variants of this method, presenting spikes only at the first exposure time (first exposure), at a time proportional to the original input intensity (latency exposed) and at all exposure time steps (direct exposure).
@@
Figure 1 outlines our complete approach to supervised RFI detection with SNNs using delta-exposure encoding as an example. Notably, the input width of the SNN either exactly matches or doubles (in the case of step-forward encodings) the number of input frequency channels ($32$ in this case). For training purposes, we slice each spectrogram into $32\times 32$ pixel patches and encode each into a spike train using the swappable encoding method. The entire spike train is fed into the SNN along the time axis, and supervision is performed on the SNN’s output spike train. At inference and final testing, the SNN’s output spike train is then decoded into another patch that is then re-stitched into a complete spectrogram for accuracy testing. This task is challenging for an SNN as we present each original time step in the spectrogram precisely once; identifying features that span the entire spike train requires the network to retain some of this information in the leaky dynamics of the neurons themselves; these results represent a baseline level of performance given what is as lightweight a network as possible for this task.
@@
Figure 1: SNN-Based RFI Detection Workflow. (a) Encoding: A spectrogram is divided into smaller patches and encoded into spikes using methods like delta-exposure encoding (shown with an exposure time of four). (b) SNN Processing: The encoded spikes are input into an SNN with a structure matching the spectrogram’s frequency channels. The SNN identifies RFI patterns from the time-varying data. (c) Decoding: Output spikes are decoded to produce an RFI mask for each patch, which is reassembled into a complete spectrogram RFI mask.
@@
(a) Original LOFAR spectrogram with latency encoding (four time step exposure). RFI features are difficult to distinguish from background information.
(b) Spectrogram after divisive normalisation. The background gradient is reduced, enhancing contrast and making RFI features more prominent, like the three bars on the right.
Figure 2: Effects of divisive normalisation on spectrogram pre-processing for RFI detection.
@@
In addition to experimenting with different encoding methods, we wanted to experiment with network sizes and neuron parameters through hyper-parameter optimisation. The HERA dataset, while artificial, contains simulated RFI sources that reflect real-world sources. Full details of this dataset are available in the methods section. For all trials, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), the Area Under the Precision-Recall Curve (AUPRC), per-pixel Accuracy, and the F1-Score. The definitions and significance of these values are found in the methods section. Owing to the relatively low per-pixel prevalence of RFI in both datasets, we expect high accuracy and AUROC values but pay greater attention to AUPRC and F1-Score, class-balanced performance measures. Our previous investigation into encoding methods tuned a restricted set of parameters, fixing the network architecture but varying batch size and training epochs. We subsequently validated the parameters discovered with repeat trials and presented the results at an earlier conference venue [25]. These results with additional results containing replicated experiments with the new delta-exposure encoding are available in the supplementary information section, but we will summarise our findings below. We found our SNN approach to outperform equivalent ANN networks, and informed our decision to focus on delta-exposure, latency and step-forward direct encodings in the new work presented in this article.
@@
Table 1: Hyper-parameter search using Optuna multi-variate optimisation for the synthetic HERA dataset.
Encoding Method Accuracy AUROC AUPRC F1-Score Num Hidden Num Layers Beta Exposure
@@
00footnotetext: We show the trials that performed best in at least one metric per encoding method. Best scores are in bold. Divisive Normalisation abbreviated to ‘DN’. ‘Num Hidden’ refers to the number of neurons in each hidden layer, and ‘Num Layers’ refers to the number of hidden layers. The best overall encoding method is underlined. Each entry scored the highest in at least one performance measure, hence the variable number of entries per encoding method.
@@
Table 2: The final results show the performance of each encoding method using the final hyper-parameters on the HERA dataset.
@@
00footnotetext: Each metric is listed as mean and standard deviation. Ten trials were completed for each encoding method. Divisive Normalisation abbreviated to ‘DN’. The best scores are bolded. Latency encoding with divisive normalisation offers superior performance in two metrics. Divisive normalisation significantly improves accuracy, AUPRC, and F1 scores for all methods while degrading AUROC scores slightly for the delta-exposure and step-forward-direct encodings.
Figure 3 presents an example spectrogram original and divisible normalised, latency encoded inference, and the original ground-truth RFI mask. This figure clearly shows that divisive normalisation removes almost all of the background ‘gradient’ information representing the simulated astronomical background from the dataset, resulting in significantly clearer inputs and outputs from the SNN. On average, divisive normalisation improved accuracy by 0.102, AUROC by 0.029, AUPRC by 0.108 and F1-Score by 0.168 across each spike-encoding method. We find latency encoding with divisive normalisation to be the strongest performing encoding method with an AUROC of 0.968, AUPRC of 0.914 and F1-Score of 0.907 to be comparable with state-of-the-art traditional algorithms, ANN methods and ANN2SNN conversions on this dataset. Specifically, the best-reported performance using AOFlagger [8], an operational RFI flagging scheme achieves an AUROC of 0.973, AUPRC of 0.880 and F1-Score of 0.873 [14]. The best reported ANN method [15] achieves an AUROC of 0.994, AUPRC of 0.959 and F1-Score of 0.945. Finally, the best previous SNN method, Spiking-Nearest-Latent-Neighbours (SNLN) [20] achieves an AUROC of 0.944, AUPRC of 0.920 and F1-Score of 0.953. SNLN, while performant, is computationally intensive, using a much larger network architecture and long inference times. We, therefore, claim that our method is near state-of-the-art despite significantly lower computational requirements than other comparable methods.
(a) Original spectrogram.
(b) Spectrogram after divisive normalisation, with background gradients reduced.
(c) Latency-based inference on the original spectrogram.
(d) Latency-based inference on the normalised spectrogram shows clearer RFI features.
(e) Ground-truth RFI mask.
Figure 3: Impact of Divisive Normalisation on RFI Detection in HERA Spectrograms. Divisive normalisation significantly reduces background noise while preserving key RFI features, improving inference performance of the SNN.
@@
Table 3: Hyper-parameter search using Optuna multi-variate optimisation for the real LOFAR dataset.
Encoding Method Accuracy AUROC AUPRC F1-Score Batch Size Epochs Beta Exposure
@@
00footnotetext: We show the trials that performed best in at least one metric per encoding method. The best scores are in bold. Divisive Normalisation abbreviated to ‘DN’. ‘Num Hidden’ refers to the number of neurons in each hidden layer, and ‘Num Layers’ refers to the number of hidden layers. The best overall encoding method is underlined. Each entry scored the highest in at least one performance measure, hence the variable number of entries per encoding method.
@@
In line with our approach to the HERA dataset, we took the best-performing parameters for each method and performed repeat trials to verify performance. Table 4 contains the results of these trials. We see that, on aggregate, our SNN-encoded methods perform better than the equivalently sized ANN methods. The best-performing method overall is Step-forward encoding with direct exposure; however, the significantly better AUROC score from latency encoding with divisive normalisation is noteworthy.
Table 4: The final results show the performance of each encoding method using the final hyper-parameters on the LOFAR dataset.
@@
Delta Exposure 0.973 0.007 0.149 0.009 0.556 0.002 0.120 0.017
Delta Exposure + DN 0.961 0.005 0.143 0.004 0.555 0.003 0.097 0.006
Latency 0.972 0.017 0.190 0.037 0.566 0.004 0.149 0.042
Latency + DN 0.978 0.003 0.630 0.003 0.254 0.011 0.198 0.016
@@
00footnotetext: Each metric is listed as mean and standard deviation. The best scores are bolded. Ten trials were completed for each encoding method. Divisive Normalisation abbreviated to ‘DN’. Step-forward-direct encoding without divisive normalisation offers the best overall performance in accuracy, AUPRC, and F1 scores. Latency encoding with divisive normalisation offers significantly higher AUROC performance than all other methods.
In Figure 4, the SNN outputs are noisy with or without divisive normalisation, although the inference trained on normalised data shows the RFI features more prominently. Interestingly, there appears to be subtle banding in the outputs of Figures 4(c) (horizontal) and 4(d) (vertical), the difference in orientation suggests perhaps a tendency for the networks to present a bias towards frequency isolated or time-isolated RFI respectively. While promising, the noisy output of the SNN is likely contributing to a degraded performance. On average, divisive normalisation degraded accuracy by 0.015, AUPRC by 0.109 and F1-Score by 0.021 but improved AUROC by 0.130 across each spike-encoding method. We find Step-forward-direct encoding to give the strongest performance with an AUROC of 0.262, AUPRC of 0.628, and F1-Score of 0.217. These results are compelling for such relatively lightweight architectures, however there exists a gap to the state-of-the-art. The best reported AOFlagger [8] results [14] give an AUROC of 0.788, AUPRC of 0.572 and F1-Score of 0.570. The best reported F1-score for an ANN method on this dataset is 0.742 [17]. The best reported AUROC and AUPRC values for an ANN method, Remove First Detect Later (RFDL), on this dataset are 0.989 and 0.748, respectively, and the best reported F1-score for the RFDL method is 0.660 [15]. Finally, the best ANN2SNN conversion achieves an AUROC of 0.609, AUPRC of 0.321 and F1-Score of 0.408 [20]. We additionally note that our best AUROC score from latency encoding with divisive normalisation achieves an AUROC of 0.630. These results indicate there exists potential for our time-series segmentation approach applied to real astronomical data, however some additional developments are needed to close the gap to the state-of-the-art.
(a) Original spectrogram.
(b) Spectrogram after divisive normalisation.
(c) Latency-based inference on the original spectrogram.
(d) Latency-based inference on the normalised spectrogram shows clearer RFI detection.
(e) Expert-labelled RFI mask.
Figure 4: Impact of Divisive Normalisation on RFI Detection in LOFAR Spectrograms. While divisive normalisation improves clarity, the SNN output still exhibits some noise compared to expert labelling.

### v2

@@ Abstract
Spiking Neural Networks (SNNs) promise efficient and dynamic spatio-temporal data processing. This paper reformulates a significant challenge in radio astronomy, Radio Frequency Interference (RFI) detection, as a time-series segmentation task suited for SNN execution. Automated systems capable of real-time operation with minimal energy consumption are increasingly important in modern radio telescopes. We explore several spectrogram encoding methods and network parameters, applying first and second-order leaky integrate and fire SNNs to tackle RFI detection. We introduce a divisive normalisation-inspired pre-processing step, improving detection performance across multiple encodings strategies. Our approach achieves competitive performance on a synthetic dataset and compelling initial results on real data from the Low-Frequency Array (LOFAR). We position SNNs as a viable path towards real-time RFI detection, with many possibilities for follow-up studies. These findings highlight the potential for SNNs performing complex time-series tasks, paving the way towards efficient, real-time processing in radio astronomy and other data-intensive fields.
@@
Introduction
Spiking Neural Networks (SNNs) are inspired by how biological neurons communicate. Instead of using continuous signals, as in Artificial Neural Networks (ANNs), they transmit information in discrete spikes or pulses, closer to how natural neurons fire [1]. ANNs have proven to be highly effective, especially when deployed on conventional computing hardware platforms, and they excel in numerous tasks when combined with large datasets. The central question remains whether SNNs can reliably achieve comparable or superior results by mimicking biological neurons more closely while using less energy and computational power, particularly when running on neuromorphic computing systems [2]. The potential for SNNs to outperform traditional models hinges on developing efficient neuromorphic hardware and identifying data-intensive applications that can exploit the temporal dynamics unique to SNNs [3, 4]. While SNNs have shown promise in several domains, particularly those leveraging neuromorphic sensing hardware, tackling increasingly demanding tasks is vital to moving the field forward.
@@ Introduction
Both conventional and machine-learned approaches treat RFI detection as a two-dimensional semantic segmentation or anomaly detection problem, where spectrograms, or ‘visibilities,’ are treated like images. Although effective, existing two-dimensional approaches to RFI detection necessitate completing an entire observation before processing can commence. This is a significant limitation as data destined to be removed is stored and processed in costly high-performance computing facilities [5]. As such, there is a strong imperative in the field to develop RFI detection techniques that are accurate and capable of operating in near real-time, with minimal energy consumption, and where applicable, minimal training data [15, 18]. Previous works have investigated real-time RFI detection surveying common mitigation strategies at various observatories [19], developing scalable real-time flaggers for large-scale arrays such as LOFAR [20] and the Apertif Radio Transient System (ARTS) at the Westerbork Synthesis Radio Telescope (WSRT) [21], and investigating the precise effects of persistent but still dynamic satellite sources and their effects on precision science experiments [22].
@@
Previous applications of SNNs to spectrographic data, such as audio processing, have primarily focused on classification tasks [26, 27]. In such tasks, a time-varying input is processed, but the final output is reduced to a single value or decision derived from the final layer of the network. However, the problem of RFI detection requires the output to retain the same temporal and spatial resolution as the input data. In this work, we reformulate RFI detection as a time-series segmentation problem, where the network must output a prediction that spans the same time and frequency dimensions as the input spectrogram to explore how well first and second-order leaky-integrate and fire (LiF) spiking neurons can capture temporal information in this task. The comparison between neuron types is also significant, as second order neurons are often more difficult to train. We report on the best-performing neuron type for each dataset in the main-text, and include the remaining results in Supplementary Notes 1 and 2. To our knowledge, this is the first time SNNs have been applied to this kind of task, and until now, it has been unknown whether backpropagation through time (BPTT) can handle the complexity of such tasks.
@@
This paper provides a robust formulation of RFI as a time-series segmentation task appropriate for direct SNN execution, investigates encoding visibility data with latency, rate, delta-modulation, delta-exposure and three distinct variations on step-forward encoding, extensive hyper-parameter optimisation, a divisive normalisation inspired pre-processing step which greatly improves detection performance across several encoding methods, and results for SNNs trained and optimised with each encoding for a synthetic dataset simulating the Hydrogen Epoch of Reionisation Array (HERA) in South Africa, and a real RFI detection dataset derived from the Low-Frequency Array (LOFAR) instrument in the Netherlands with comparison to identically sized ANN feed-forward networks. We opt to compare results to directly sized, very simple ANNs to highlight the specific differences introduced by using spiking neurons specifically. This paper builds upon our previously presented work [30], extending it in several important ways. The delta-exposure encoding method and divisive-normalisation-inspired pre-processing step are previously unseen techniques. We provide more comprehensive hyper-parameter optimisation extending into network depth and width in addition to a more extensive analysis of the resulting parameter space. Lastly, by including results for the real LOFAR dataset, we present the first study, to the best of our knowledge, demonstrating the application of from-scratch trained SNNs on real radio astronomy data. In doing so, we achieve state-of-the-art RFI detection performance on the synthetic HERA dataset and initial results on the real LOFAR dataset, demonstrating the efficacy of our approach while employing small multi-layer feed-forward SNNs. This investigation demonstrates the capability of SNNs to perform a data-intensive complex time-series segmentation task in the form of RFI detection with relatively simple network architectures, paving the way for more sophisticated SNN-based RFI detection schemes and critically finding a scientifically impactful problem in which SNNs and neuromorphic computing may hold unique advantages over contemporary methods.
@@ Results
RFI Detection as Time-Series Segmentation
@@
This formulation tasks a classifier to segment the time series, in this case, as boolean values rather than a two-dimensional image. In the case of an ANN, the encoding and decoding functions are the identity. This formulation leaves encoding and decoding visibility data as spikes for this specific task open for experimentation. The spike encoding process involves stretching the original time steps in the spectrograms into longer exposures, allowing the SNN to react to the current input. We experimented with seven encoding methods. The methods section presents detailed formulations and examples of each encoding approach.
@@
Figure 1 outlines our complete approach to supervised RFI detection with SNNs using delta-exposure encoding as an example. Notably, the input width of the SNN either exactly matches or doubles (in the case of step-forward encodings) the number of input frequency channels ($32$ in this case). For training purposes, we slice each spectrogram into $32\times 32$ pixel patches and encode each into a spike train using the swappable encoding method. Without patching, training an SNN with BPTT becomes very challenging as the memory requirements grows proportional to the length of the input spike-train. We discuss difficulties introduced by patching the inputs in the Methods section. Bounding the number of time-steps allows for tractable learning as the network only needs to learn coherent behaviour for a limited amount of time. The entire spike train is fed into the SNN along the time axis, and supervision is performed on the SNN’s output spike train. At inference and final testing, the SNN’s output spike train is then decoded into another patch that is then re-stitched into a complete spectrogram for accuracy testing. This task is challenging for an SNN as we present each original time step in the spectrogram precisely once; identifying features that span the entire spike train requires the network to retain some of this information in the leaky dynamics of the neurons themselves; these results represent a baseline level of performance given what is as lightweight a network as possible for this task.
@@
Figure 1: Spiking Neural Network (SNN) Based Radio Frequency Interference (RFI) Detection Workflow. (a) Encoding: A spectrogram is divided into smaller patches and encoded into spikes using methods like delta-exposure encoding (shown with an exposure time of four). (b) SNN Processing: The encoded spikes are input into an SNN with a structure matching the spectrogram’s frequency channels. The SNN identifies RFI patterns from the time-varying data. (c) Decoding: Output spikes are decoded to produce an RFI mask for each patch, which is reassembled into a complete spectrogram RFI mask.
@@
Figure 2: Effects of divisive normalisation on spectrogram pre-processing for Radio Frequency Interference (RFI) detection. Panel (a) depicts an original full LOFAR spectrogram (left) and sub-section patch (right) encoded with latency encoding spike raster (bottom, four time step exposure). RFI features are difficult to distinguish from background information. Panel (b) depicts the same full spectrogram (left), sub-section patch (right) after divisive normalisation with a latency encoding spike raster (bottom, four time step exposure). The background gradient is reduced, increasing spike sparsity and making RFI features more prominent, like the three bars on the right.
@@
In addition to experimenting with different encoding methods, we wanted to experiment with network sizes and neuron parameters through hyper-parameter optimisation. The HERA dataset, while artificial, contains simulated RFI sources that reflect real-world sources. Full details of this dataset are available in the methods section. For all trials, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), the Area Under the Precision-Recall Curve (AUPRC), per-pixel Accuracy, and the F1-Score. The definitions and significance of these values are found in the methods section. Owing to the relatively low per-pixel prevalence of RFI in both datasets, we expect high accuracy and AUROC values but pay greater attention to AUPRC and F1-Score, class-balanced performance measures. Our previous investigation into encoding methods tuned a restricted set of parameters, fixing the network architecture but varying batch size and training epochs. We subsequently validated the parameters discovered with repeat trials and presented the results at an earlier conference venue [30]. We summarise our findings of this previous work below. We found our SNN approach to outperform equivalent ANN networks, and informed our decision to focus on delta-exposure, latency and step-forward direct encodings in the new work presented in this article.
@@
Table 1: Hyper-parameter search using Optuna multi-variate optimisation for the synthetic HERA dataset.00footnotetext: We show the trials that performed best in at least one metric per encoding method. Best scores are in bold. Abbreviations: Divisive Normalisation ‘DN’; Artificial Neural Network ‘ANN’; Area Under the Receiver Operator Curve ‘AUROC’, and Area Under the Precision Recall Curve ‘AUPRC’. ‘Num Hidden’ refers to the number of neurons in each hidden layer, and ‘Num Layers’ refers to the number of hidden layers. The best overall encoding method is underlined. Each entry scored the highest in at least one performance measure, hence the variable number of entries per encoding method.
Encoding Method Accuracy AUROC AUPRC F1-Score
Num
Hidden
Num
Layers
Beta Exposure
@@
Table 2: The final results show the performance of each encoding method using the final hyper-parameters on the HERA dataset with first-order LiF neurons.00footnotetext: Each metric is listed as mean and standard deviation. Ten trials were completed for each encoding method. Abbreviations: Divisive Normalisation ‘DN’; Artificial Neural Network ‘ANN’; Area Under the Receiver Operator Curve ‘AUROC’, and Area Under the Precision Recall Curve ‘AUPRC’. The best scores are bolded. Latency encoding with divisive normalisation offers superior performance in two metrics. Divisive normalisation significantly improves accuracy, AUPRC, and F1 scores for all methods while degrading AUROC scores slightly for the delta-exposure and step-forward-direct encodings.
@@
Figure 3 presents an example spectrogram original and divisible normalised, latency encoded inference, and the original ground-truth RFI mask. This figure clearly shows that divisive normalisation removes almost all of the background ‘gradient’ information representing the simulated astronomical background from the dataset, resulting in significantly clearer inputs and outputs from the SNN. Interestingly, we find using a more complex second-order LiF neuron model degrades performance significantly on this dataset and provide these results in Supplementary Table 1. On average, divisive normalisation improved accuracy by 0.102, AUROC by 0.029, AUPRC by 0.108 and F1-Score by 0.168 across each spike-encoding method. We find latency encoding with divisive normalisation to be the strongest performing encoding method with an AUROC of 0.968, AUPRC of 0.914 and F1-Score of 0.907 to be comparable with state-of-the-art traditional algorithms, ANN methods and ANN2SNN conversions on this dataset. Table 3 contains a comparison between this work and the current state of the art algorithmic, ANN and SNN performances on a HERA-derived dataset. We see that despite the drastically reduced model size, our directly trained BPTT + DN latency encoded model performs competitively against existing methods.
Table 3: Comparison to other algorithmic, ANN and SNN baselines on the HERA dataset.00footnotetext: Abbreviations: Divisive Normalisation ‘DN’; Spiking Neural Network ‘SNN’; Artificial Neural Network ‘ANN’; Backpropagation Through Time ‘BPTT’; Area Under the Receiver Operator Curve ‘AUROC’, and Area Under the Precision Recall Curve ‘AUPRC’. We see that directly trained SNNs perform comparably to the state of the art, which is noteworthy considering their relatively small size.
Work Model Method AUROC AUPRC F1
[15] AOFlagger Algorithm 0.974 0.880 0.873
[16] Auto-Encoder ANN 0.994 0.959 0.945
[25]
Auto-Encoder
ANN2SNN 0.944 0.920 0.953
This work BPTT + DN SNN 0.968 0.914 0.907
Figure 3: Impact of Divisive Normalisation on Radio Frequency Inference (RFI) Detection in HERA Spectrograms. Sub-panels depict: (a) original spectrogram; (b) spectrogram after divisive normalisation, with background gradients reduced; (c) latency-based inference on the original spectrogram; (d) latency-based inference on the normalised spectrogram showing clearer RFI features; (e) residual between original inference and mask; (f) residual between normalised inference and mask, and (g) ground-truth RFI mask. Divisive normalisation significantly reduces background noise while preserving key RFI features, improving inference performance of the SNN.
@@
Table 4: Hyper-parameter search using Optuna multi-variate optimisation for the real LOFAR dataset.00footnotetext: We show the trials that performed best in at least one metric per encoding method. The best scores are in bold. Abbreviations: Divisive Normalisation ‘DN’; Artificial Neural Network ‘ANN’; Area Under the Receiver Operator Curve ‘AUROC’, and Area Under the Precision Recall Curve ‘AUPRC’. ‘Num Hidden’ refers to the number of neurons in each hidden layer, and ‘Num Layers’ refers to the number of hidden layers. The best overall encoding method is underlined. Each entry scored the highest in at least one performance measure, hence the variable number of entries per encoding method.
Encoding Method Accuracy AUROC AUPRC F1-Score
Num
Hidden
Num
Layers
Beta Exposure
@@
In line with our approach to the HERA dataset, we took the best-performing parameters for each method and performed repeat trials to verify performance. Table 2 contains the results of these trials. We see that, on aggregate, our SNN-encoded methods perform better than the equivalently sized ANN methods. The best-performing method overall is latency encoding; however, the significantly better AUPRC score from step-forward (direct) encoding with divisive normalisation is noteworthy.
Table 5: The final results show the performance of each encoding method using the final hyper-parameters on the LOFAR dataset with second-order LiF neurons.00footnotetext: Each metric is listed as mean and standard deviation. The best scores are bolded. Ten trials were completed for each encoding method. Abbreviations: Divisive Normalisation ‘DN’; Artificial Neural Network ‘ANN’; Area Under the Receiver Operator Curve ‘AUROC’, and Area Under the Precision Recall Curve ‘AUPRC’. Latency encoding without divisive normalisation offers the best overall performance in accuracy and F1 scores. Step-Forward (direct) encoding with divisive normalisation offers significantly higher AUPRC performance than all other methods. An identically sized ANN does perform best in AUROC however.
@@
Delta Exposure 0.887 0.294 0.314 0.007 0.621 0.108 0.397 0.009
Delta Exposure + DN 0.992 0.000 0.312 0.000 0.629 0.011 0.395 0.001
@@
Latency + DN 0.989 0.003 0.335 0.005 0.475 0.007 0.438 0.005
Step-Forward-Direct 0.989 0.006 0.351 0.008 0.497 0.015 0.469 0.012
Step-Forward-Direct + DN 0.992 0.000 0.311 0.001 0.693 0.000 0.393 0.002
@@
In Figure 4, the SNN outputs are noisy with or without divisive normalisation, although the inference trained on normalised data shows the RFI features more prominently. We also note the observation of banding in frequency and time dimensions near patch boundaries in some trials, which is the result of biases learned by these network instances and a fuzziness in the decoding process, which is the result of attempting learning on a significantly more complex dataset. We discuss this phenomena and other preprocessing techniques, which have been key to achieving competitive performance in the Methods section. Moreover, we provide significantly poorer results using the simpler first-order LiF neuron model in Supplementary Table 2. While promising, the noisy output of the SNN is likely contributing to a degraded performance. On average, divisive normalisation increased accuracy by 0.035, decreased AUROC by 0.017, increased AUPRC by 0.025 but decreased F1-Scores by 0.038 across all spike-encoding methods. We find Latency encoding to give the strongest performance with an AUROC of 0.346, AUPRC of 0.604, and F1-Score of 0.474. These results are compelling for such relatively lightweight architectures, however there exists a gap to the state-of-the-art. Table 6 compares the results of our best-performing directly trained SNN (Step-Forward with direct encoding) with existing algorithmic, ANN, and SNN baselines. We see that there is a sizeable gap to the state of the art, indicating potential for this time-series segmentation approach to apply to real-world data, however some additional developments are needed to close the gap to the state-of-the-art. Notably however, these results exceed prior ANN2SNN conversion-based results.
Table 6: Comparison to other algorithmic, ANN, and SNN baselines on the LOFAR dataset.00footnotetext: Abbreviations: Divisive Normalisation ‘DN’; Spiking Neural Network ‘SNN’; Artificial Neural Network ‘ANN’; Backpropagation Through Time ‘BPTT’; Area Under the Receiver Operator Curve ‘AUROC’, and Area Under the Precision Recall Curve ‘AUPRC’. While performance still lags the state of the art, these results encourage further development considering the compact size of our BPTT trained SNNs.
Work Model Method AUROC AUPRC F1
[15] AOFlagger Algorithm 0.788 0.572 0.570
[18] Auto-Encoder ANN 0.742
[16] Auto-Encoder ANN 0.989 0.748 0.660
[25]
Auto-Encoder
ANN2SNN 0.609 0.321 0.408
This work BPTT SNN 0.346 0.604 0.474
Figure 4: Impact of Divisive Normalisation on Radio Frequency Interferrence (RFI) Detection in LOFAR Spectrograms. Sub-panels depict: (a) original spectrogram; (b) spectrogram after divisive normalisation; (c) latency-based inference on the original spectrogram; (d) latency-based inference on the normalised spectrogram showing a significantly more conservative flagging approach; (e) residual between original inference and mask; (f) residual between normalised inference and mask, and (g) expert-labelled RFI mask. While divisive normalisation improves clarity, the SNN output still exhibits significant noise compared to expert labelling. This is most apparent in the residual plots where in both cases key features are missing from inference.